OpenStudy (anonymous):

x^5 − x^4 + 1 find upper and lower bounds

OpenStudy (lochana):

y = x^5 - x^4 + 1 dy/dx = 5x^4 - 4x^3 =x^3(5x-4) upper x = 0 lower x= 4/5

OpenStudy (anonymous):

Thanks! That's the same answer I got, but I typed it into webassign, and it graded me as wrong :/ Weird.

OpenStudy (anonymous):

The real answer was Upper:1 Lower:-1

OpenStudy (anonymous):

I figured it out! Thanks. I checked with synthetic division.